OpenStudy (anonymous):

What is the sum of a 6-term geometric series if the first term is 19 and the last term is 319,333? (3 points) A) 372,552 B) 392,160 C) 411,768 D) 431,376 What is the sum of an 8-term geometric series if the first term is 15 and the last term is -4,199,040? (3 points) A) -3,599,175 B) -3,359,230 C) -3,119,285 D) -2,879,340

OpenStudy (campbell_st):

find the common ratio use \[T _{n}=ar^{n-1}\] and the information for the 6th term \[319333 = 19r^5\] then \[16807 = r^5\] r = 7 so you know a = 19 r = 7 and n = 6 use the sum or a geometric series formula \[s _{n} = \frac{a(r^n - 1)}{r - 1}\] substitute and evaluate for your answer

OpenStudy (anonymous):

2235312 i got tht for an answer :/

OpenStudy (campbell_st):

well something is amiss...

OpenStudy (campbell_st):

are you sure there isn't a negative sign missing from 319333

OpenStudy (anonymous):

nope

OpenStudy (campbell_st):

nope that doesn't work either

OpenStudy (campbell_st):

for B I found the common ratio to be -6

OpenStudy (anonymous):

I got A for the first one.

OpenStudy (anonymous):

And A for the second, also.

OpenStudy (anonymous):

thank you so muchhh

OpenStudy (campbell_st):

for the 2nd question, a=15, n = 8 and r = -6 will give answer A for question 1 I got answer A \[s _{6}= \frac{19(7^6 -1)}{7-1}\]